# CMAR-Net: Accurate Cross-Modal 3D SAR Reconstruction of Vehicle Targets with Sparse-Aspect Multi-Baseline Data

### v4

@@ III-B Results for Simulated Data
Figure 8 presents the 3D reconstruction results of CMAR-Net under an SNR of 30 dB and with 8 sub-aspects, alongside the corresponding BP imaging results and 3D model visualizations. To comprehensively illustrate the reconstruction quality, the CMAR-Net output is visualized from three viewing angles with (azimuth, elevation) parameters of (0°, 17°), (30°, 17°), and (135°, 17°). For comparison, the BP imaging results and its corresponding 3D model are shown using the (0°, 17°) viewpoint. Unless otherwise specified, this viewpoint is used by default in all subsequent experimental visualizations. As shown in the figure, the BP method, limited by the use of only 8 sub-aspect observations, produces significant sidelobe artifacts and incomplete vehicle structures. In contrast, CMAR-Net substantially enhances the reconstruction of the vehicle surface, successfully recovering visually distinguishable geometric features. The resulting shape closely matches the corresponding 3D model.
Figure 10: Imaging results of different methods at an SNR of 25dB and 12 sub-aspects.
Figure 11: Imaging results of different methods at an SNR of 20dB and 10 sub-aspects.
Figure 12: Imaging results of different methods at an SNR of 15dB and 8 sub-aspects.
Figure 13: Imaging results of different methods at an SNR of 10dB and 6 sub-aspects.
Figure 14: Imaging results of different methods at an SNR of 5dB and 4 sub-aspects.
@@ Methods
We compare CMAR-Net with baseline methods. Figure 9 presents the imaging results of different methods at an SNR of 30dB and 8 sub-aspects. The BP method preserves the strong scattering points of the vehicle body by non-coherently integrating the maximum amplitude values from sub-aspect images. For instance, the bucket structure of Pickup are prominently improved. However, the imaging quality is degraded by high side lobes, as evidenced by the distorted front and windshield of the MPV and Camry. The CS method mitigates the impact of side lobes by optimizing sub-aspect imaging. However, the imaging results reveal significant structural omissions, attributed to the sparsity assumptions and the configuration of the optimization parameters.
DL methods supervised by radar images leverage prior knowledge from full-aperture imaging, enabling the reconstruction of vehicle images with more scattering details using sparse data. Such as the clearer bucket on the Pickup. However, due to constraints imposed by electromagnetic characteristics, the imaging results lack well-defined planar structures on the vehicle body.
To evaluate the reconstruction performance of CMAR-Net under different conditions, we conduct experiments where the SNR decreased from 25 dB to 5 dB, and the number of sub-aspects reduced from 12 to 4. Figures 10 to 14 present the imaging results for all experimental conditions.
As the SNR decreases and the number of sub-aspects reduces, the imaging results of the BP and CS methods show more side lobes, leading to reduced vehicle recognizability. SACNet and UNet3D demonstrate strong robustness under deteriorating imaging conditions, providing consistent reconstruction results. After integrating the cross-modal supervision, their performance remains stable. In contrast, CMAR-Net consistently delivers the best imaging performance across all vehicle types. Even under the most challenging conditions (5 dB SNR and 4 sub-aspects), CMAR-Net achieves excellent reconstruction, while the performance of other methods significantly deteriorates.
III-B3 Quantitative comparison
We calculate the quality metrics under different data conditions, with the results presented in Tables II to III. Figures 15 visualizes these results using curve plots. The data reveal that all existing methods perform worse across all metrics compared to DL-based methods with cross-modal supervision, with particularly notable differences in the IoU and CE metrics. This disparity is largely due to modality differences between the output images and the ground truth data. And among the proposed DL-based methods, CMAR-Net consistently achieves the best performance across all metrics. The only exception occurs under the condition of 4 sub-aspects and an SNR of 5 dB, where CMAR-Net’s SSIM is slightly lower than the method using UNet3D as the backbone. We believe this is due to the UNet3D’s tendency to output higher pixel values (closer to 1). However, this behavior in fact leads to a loss of structural information, exhibiting poor generalization of the network, a conclusion that will be confirmed in real-world experiments.
In summary, compared to traditional methods and DL methods supervised with radar images, CMAR-Net achieved improvements of $>$100% in IoU, 91.23% in CE, 80.28% in PSNR and 39.32% in SSIM. When compared to DL methods with cross-modal supervision, CMAR-Net shows enhancements of 22.54% in IoU, 52.84% in CE, 24.32% in PSNR and 13.09% in SSIM. These results demonstrate that CMAR-Net can reconstruct more realistic images than other methods.
@@ III-C1 Real-world dataset
The real-world data used in this study is sourced from the GOTCHA Circular SAR dataset, collected and released by the United States Air Force Laboratory[43]. This dataset captures a scene with various civilian vehicles using a radar with the center frequency of 9.6 GHz and the bandwidth of 640 MHz. The data is acquired in circular SAR mode through 8 passes at different altitudes, with elevation angles of [45.66°, 44.01°, 43.92°, 44.18°, 44.14°, 43.53°, 43.01°, 43.06°]. The combination of circular observation aperture and altitude diversity enabled 3D imaging of the entire scene. We select data from two SUVs and two sedans. Figure 17 illustrates the scene, radar flight passes, and vehicle photos.
@@
Figure 18: Imaging results of different methods using real-world data. Imaging using 12 sub-aspects
Figure 19: Imaging results of different methods using real-world data. Imaging using 10 sub-aspects
Figure 20: Imaging results of different methods using real-world data. Imaging using 8 sub-aspects
Figure 21: Imaging results of different methods using real-world data. Imaging using 6 sub-aspects
Figure 22: Imaging results of different methods using real-world data. Imaging using 4 sub-aspects
@@ III-C2 Imaging results
Figure 18 to 22 present the imaging results on real-world data with varying number of sub-aspects. To facilitate observation, the visualization angle is set to (0°, 0°) when presenting the results. The BP method produces results heavily affected by noise and side lobes, which can hardly be recognized. The CS method alleviates these issues, but the reconstructions still appear as chaotic point clouds, leaving vehicle features unrecognizable, especially when using highly sparse data. For example, when the number of sub-aspects is 4, traditional methods completely fail, producing totally disordered point clusters without any discernible or meaningful vehicle features.
In contrast, radar image-supervised DL methods can reconstruct vaguely discernible vehicle contours. However, as the number of aspects decreases, these contours deteriorates. For example, when the number of sub-aspects is reduced to 4, SACNet and UNet3D show significant omissions in reconstructing the roof of Jeep. After integrating cross-modal supervision, SACNet and UNet3D exhibit weak generalization when processing real-world data, struggling to accurately reconstruct the vehicle’s geometry. Particularly, when the number of sub-aspects is 4, SACNet fails completely and cannot output any visible features.
Tables IV-V and Figure 23 show the evaluation metrics. Compared to traditional methods and DL methods supervised with radar images, CMAR-Net achieved improvements of $>$100% in IoU, 89.36% in CE, 75.83% in PSNR and 47.85% in SSIM. When compared to DL methods with cross-modal supervision, CMAR-Net shows enhancements of 86.42% in IoU, 67.99% in CE, 38.81% in PSNR and 10.75% in SSIM.
Figure 23: Image quality metrics-number of sub-aspect curves for different methods on real-world data.
@@
Methods Number of sub-aspect
@@
Methods Number of sub-aspect
@@ III-C3 Ablation study
The 7th and 8th sets of data in the Figure 18-22 and table IV-V show the imaging results and evaluation metrics after ablating the PRP unit and data augmentation strategy. Without the PRP unit, the vehicle images become blurry and suffer greater pixel loss as the number of aspects decreases. And in the absence of data augmentation, the network consistently reconstructs all vehicle type as sedan type, revealing severe overfitting. The PRP unit design improves SSIM, PSNR, IoU, and CE by 8.25%, 21.25%, 24.15%, and 42.82%, respectively. Data augmentation contributes improvements of 8.65%, 19.92%, 42.94%, and 46.75%. These results clearly demonstrate the effectiveness of both the PRP unit and data augmentation strategy.
@@ IV Conclusion
In this paper, we propose a DL-based multi-baseline SAR 3D reconstruction method that utilizes only sparse aspect data. We introduce the concept of cross-modal learning and design a DNN tailored for 3D SAR reconstruction—CMAR-Net. By leveraging differentiable rendering, CMAR-Net can be effectively trained, addressing the key challenge of supervising 3D reconstruction network with 2D optical images. Extensive experiments on both simulated and real-world datasets show that CMAR-Net significantly outperforms all existing baseline methods, including traditional imaging algorithms and DL-based approaches, in terms of image quality and generalization. Moreover, the existing DL-based SAR reconstruction methods, adopting our proposed cross-modal learning paradigm, experience a significant performance improvement, which highlights the superiority and versatility. Last but not least, our method eliminates the need for time-consuming preprocessing of full-aperture data. Instead, it only needs 2D images obtained through computer rendering or camera capture. This significantly reduces the difficulty of dataset construction, greatly enhancing the practicality of 3D SAR reconstruction technology.
Our approach offers new insights for 3D SAR researchers and is expected to comprehensively improve the research landscape of 3D SAR imaging. The technique has the potential to be applied to fields such as aerial remote sensing, terrain mapping, and obstacle detection. In the future, researchers can expand the scope to 3D reconstruction of other scenes (e.g., buildings, streets), or explore fusion with other heterogeneous data (e.g., infrared data), and so on.

### v6

@@ III-B Results for Simulated Data
Figure 9: Overview of 3D reconstruction performance of CMAR-Net. BP results (1st row), CMAR-Net reconstructions (2nd-4th rows) and ground truth (5th row). Imaging at an SNR of 30dB and 8 observation aspects.
Figure 10: Imaging results of different methods. Each column corresponds to an imaging method, and each row represents a vehicle type. Imaging at an SNR of 30dB and 8 observation aspects.
Figure 9 shows the 3D reconstruction results of CMAR-Net under an SNR of 30 dB with 8 observation aspects, together with the corresponding BP imaging results and 3D model visualizations. To better illustrate reconstruction quality, the CMAR-Net output is rendered from three viewpoints, specified by (azimuth, elevation) = (0°, 17°), (30°, 17°), and (135°, 17°). For comparison, the BP result and its corresponding 3D model are visualized at the (0°, 17°) viewpoint, which is also used as the default viewpoint in all subsequent experimental results unless otherwise noted. As shown, the BP method—constrained by the limited number of observation aspects—produces strong sidelobe artifacts and incomplete vehicle structures. In contrast, CMAR-Net markedly improves the reconstruction of the vehicle surface, recovering visually distinguishable geometric features, and the resulting shape closely matches the reference 3D model.
@@ III-B2 Comparison with Baseline Methods
Figure 11: Imaging results of different methods at an SNR of 25dB and 12 observation aspects.
Figure 12: Imaging results of different methods at an SNR of 20dB and 10 observation aspects.
Figure 13: Imaging results of different methods at an SNR of 15dB and 8 observation aspects.
Figure 14: Imaging results of different methods at an SNR of 10dB and 6 observation aspects.
Figure 15: Imaging results of different methods at an SNR of 5dB and 4 observation aspects.
We compare CMAR-Net with baseline methods. Figure 10 presents the imaging results of different methods at an SNR of 30dB and 8 observation aspects. The BP method preserves the strong scattering points of the vehicle body by non-coherently integrating the maximum amplitude values from sub-aspect images. For instance, the bucket structure of Pickup are prominently improved. However, the imaging quality is degraded by high side lobes, as evidenced by the distorted front and windshield of the MPV and Camry. The CS method mitigates the impact of side lobes by optimizing sub-aspect imaging. However, the imaging results reveal significant structural omissions, attributed to the sparsity assumptions and the configuration of the optimization parameters.
DL methods supervised by radar images leverage prior knowledge from full-aspect (360°) imaging, enabling the reconstruction of vehicle images with more scattering details using sparse data. Such as the clearer bucket on the Pickup. However, due to constraints imposed by electromagnetic characteristics, the imaging results lack well-defined planar structures on the vehicle body.
To evaluate the reconstruction performance of CMAR-Net under different conditions, we conduct experiments where the SNR decreased from 25 dB to 5 dB, and the number of observation aspects reduced from 12 to 4. Figures 11 to 15 present the imaging results for all experimental conditions.
As the SNR decreases and the number of observation aspects reduces, the imaging results of the BP and CS methods show more side lobes, leading to reduced vehicle recognizability. SACNet and UNet3D demonstrate strong robustness under deteriorating imaging conditions, providing consistent reconstruction results. After integrating the cross-modal supervision, their performance remains stable. In contrast, CMAR-Net consistently delivers the best imaging performance across all vehicle types. Even under the most challenging conditions (5 dB SNR and 4 observation aspects), CMAR-Net achieves excellent reconstruction, while the performance of other methods significantly deteriorates.
III-B3 Quantitative comparison
We calculate the quality metrics under different data conditions, with the results presented in Tables II to III. Figures 16 visualizes these results using curve plots. The data reveal that all existing methods perform worse across all metrics compared to DL-based methods with cross-modal supervision, with particularly notable differences in the IoU and CE metrics. This disparity is largely due to modality differences between the output images and the ground truth data. And among the proposed DL-based methods, CMAR-Net consistently achieves the best performance across all metrics. The only exception occurs under the condition of 4 aspects and an SNR of 5 dB, where CMAR-Net’s SSIM is slightly lower than the method using UNet3D as the backbone. We believe this is due to the UNet3D’s tendency to output higher pixel values (closer to 1). However, this behavior in fact leads to a loss of structural information, exhibiting poor generalization of the network, a conclusion that will be confirmed in real-world experiments.
In summary, compared to traditional methods and DL methods supervised with radar images, CMAR-Net achieved improvements of >>100% in IoU, 91.23% in CE, 80.28% in PSNR and 39.32% in SSIM. When compared to DL methods with cross-modal supervision, CMAR-Net shows enhancements of 22.54% in IoU, 52.84% in CE, 24.32% in PSNR and 13.09% in SSIM. These results demonstrate that CMAR-Net can reconstruct more realistic images than other methods.
@@ III-C1 Real-world dataset
The real-world data used in this study is sourced from the GOTCHA Circular SAR dataset, collected and released by the United States Air Force Laboratory[16]. This dataset captures a scene with various civilian vehicles using a radar with the center frequency of 9.6 GHz and the bandwidth of 640 MHz. The data is acquired in circular SAR mode through 8 baselines at different altitudes, with elevation angles of [45.66°, 44.01°, 43.92°, 44.18°, 44.14°, 43.53°, 43.01°, 43.06°]. The combination of circular observation aspect and altitude diversity enabled 3D imaging of the entire scene. We select data from two SUVs and two sedans. Figure 18 illustrates the scene, radar flight baselines, and vehicle photos.
@@ III-C2 Imaging results
Figure 19: Imaging results of different methods using real-world data. Imaging using 12 observation aspects
Figure 20: Imaging results of different methods using real-world data. Imaging using 10 observation aspects
Figure 21: Imaging results of different methods using real-world data. Imaging using 8 observation aspects
Figure 22: Imaging results of different methods using real-world data. Imaging using 6 observation aspects
Figure 23: Imaging results of different methods using real-world data. Imaging using 4 observation aspects
Figure 19 to 23 present the imaging results on real-world data with varying number of observation aspects. To facilitate observation, the visualization angle is set to (0°, 0°) when presenting the results. The BP method produces results heavily affected by noise and side lobes, which can hardly be recognized. The CS method alleviates these issues, but the reconstructions still appear as chaotic point clouds, leaving vehicle features unrecognizable, especially when using highly sparse data. For example, when the number of observation aspects is 4, traditional methods completely fail, producing totally disordered point clusters without any discernible or meaningful vehicle features.
In contrast, radar image-supervised DL methods can reconstruct vaguely discernible vehicle contours. However, as the number of aspects decreases, these contours deteriorates. For example, when the number of aspects is reduced to 4, SACNet and UNet3D show significant omissions in reconstructing the roof of Jeep. After integrating cross-modal supervision, SACNet and UNet3D exhibit weak generalization when processing real-world data, struggling to accurately reconstruct the vehicle’s geometry. Particularly, when the number of aspects is 4, SACNet fails completely and cannot output any visible features.
Tables IV-V and Figure 24 show the evaluation metrics. Compared to traditional methods and DL methods supervised with radar images, CMAR-Net achieved improvements of >>100% in IoU, 89.36% in CE, 75.83% in PSNR and 47.85% in SSIM. When compared to DL methods with cross-modal supervision, CMAR-Net shows enhancements of 86.42% in IoU, 67.99% in CE, 38.81% in PSNR and 10.75% in SSIM.
@@
Methods Number of observation aspects
@@
Methods Number of observation aspects
@@
Figure 24: Image quality metrics-number of observation aspects curves for different methods on real-world data.
@@ III-C3 Ablation study
The 7th and 8th sets of data in the Figure 19-23 and table IV-V show the imaging results and evaluation metrics after ablating the PRP unit and data augmentation strategy. Without the PRP unit, the vehicle images become blurry and suffer greater pixel loss as the number of aspects decreases. And in the absence of data augmentation, the network consistently reconstructs all vehicle type as sedan type, revealing severe overfitting. The PRP unit design improves SSIM, PSNR, IoU, and CE by 8.25%, 21.25%, 24.15%, and 42.82%, respectively. Data augmentation contributes improvements of 8.65%, 19.92%, 42.94%, and 46.75%. These results clearly demonstrate the effectiveness of both the PRP unit and data augmentation strategy.
@@ IV Conclusion
In this paper, we propose a DL-based sparse 3D SAR reconstruction method for vehicle target. We introduce the concept of cross-modal learning and design a DNN tailored for 3D SAR reconstruction—CMAR-Net. By leveraging differentiable rendering, CMAR-Net can be effectively trained, addressing the key challenge of supervising 3D reconstruction network with 2D optical images. Extensive experiments on both simulated and real-world datasets show that CMAR-Net significantly outperforms all existing baseline methods, including traditional imaging algorithms and DL-based approaches, in terms of image quality and generalization. Moreover, the existing DL-based SAR reconstruction methods, adopting our proposed cross-modal learning paradigm, experience a significant performance improvement, which highlights the superiority and versatility. Last but not least, our method eliminates the need for time-consuming preprocessing of full-aspect data. Instead, it only needs 2D images obtained through computer rendering or camera capture. This significantly reduces the difficulty of dataset construction, greatly enhancing the practicality of DL-based 3D SAR reconstruction technology.
Our approach provides new insights into 3D SAR research, particularly advancing target imaging. By addressing key challenges in imaging quality, it can support improved applications of 3D SAR in aerial remote sensing, target interpretation, and urban surveillance. Future research could extend this framework to other complex targets, such as buildings or urban environments, and explore integration with additional modalities, such as infrared or hyperspectral data, to further enhance reconstruction performance.